# Trustworthy AI on Safety, Bias, and Privacy: A Survey

### v1

@@ 3.3 Limitations of Existing Techniques
All existing studies that aim to overcome spurious correlations work with the assumption that environments/groups that are overrepresented are the groups that contribute to the learning of spurious correlations. Reliance on this assumption makes it easy to identify which samples contain the spurious features causing problems, which allows for further representational alignment or changes to the training distribution. Mulchandani and Kim (2025) show that this assumption does not always hold in practice and that minority groups can contain spurious features that can mislead a network significantly.
@@
Works by Ahmed et al. (2021); Zhang et al. (2022) align representations of samples belonging to different environments within the same class. While effective at overcoming spurious correlations, these techniques reduce overall testing accuracies due to the loss of representational richness.
@@
Most works depend heavily on hyperparameter tuning, where they optimize for the best worst-group accuracy. Optimization is done with the help of a validation split that mimics the distribution of shifted testing environments. Such access to a validation split that mimics test-time distribution is unrealistic. Gulrajani and Lopez-Paz (2021) show that without access to such a validation set, standard Empirical Risk Minimization outperforms seemingly promising solutions.
@@ 3.4 Creating Robust Solutions: Next Steps
Most standard and state-of-the-art techniques assume an equal contribution to the learning and reliance of spurious correlations. In other words, every training sample belonging to the environment known to cause reliance on spurious correlations is treated the same way. Mulchandani and Kim (2025) show that samples within an environment contribute differently to learning of spurious correlations and show that these differences are extreme in practice. They train a network to learn gender classification, where a fraction of the male samples contain eyeglasses. In their work, the degree of spurious feature reliance is measured by observing the test accuracy of female samples containing eyeglasses (Worst-Group Accuracy). They observe that removing 97% of easy-to-understand male samples with eyeglasses has almost no improvement on the testing accuracy of female samples with eyeglasses. However, removing 10% of hard-to-understand male samples with eyeglasses doubles the testing accuracy of female samples with eyeglasses, as shown in Fig. 1. They show that such pruning has minimal impact on testing accuracy of the male class.
@@ 3.5 Intersection of Spurious Correlations with Other Areas of Study
Recent work has shown that deep neural networks have a tendency to rely on short-cut solutions or heuristics when learning to solve reasoning tasks, instead of robust rules that actually cover the solution to the problem Zhang et al. (2023); Nikankin et al. (2025). This makes it difficult for networks to generalize to different or more challenging domains. A good example of this is the length generalization problem, where a network is unable to solve simple arithmetic operations on numbers of length different from those observed during training, despite these operations requiring the same set of rules Zhou et al. (2024b); Lee et al. (2024).
@@
Yang et al. (2022) show that neural networks pick up on spurious features present in only a handful of training samples, which can lead to privacy leaks.
@@ 4.1 Privacy Correlation on Model Capacity and Data Complexity
The membership privacy risks of machine learning models are mainly caused by the model’s memorization of the training data points. This means that over-memorization is one of the sources of privacy risks Yeom et al. (2020). Carlini et al. (2022) claimed some data points must be more privacy-risky after the removal of original privacy-risky data points and retraining from scratch. This is mainly due to the relative changes between the model capacity and the data complexity. Tan et al. (2022) found that excess model capacity (a.k.a., overparameterization) is another factor of privacy risks. Additionally, Tan et al. (2023a) showed the larger-capacity model not only memorizes more on training data points than smaller networks but also memorizes faster (i.e., within fewer iterations). In fact, changing data complexity can also change the model’s memorization behavior. As shown in Fig. 2, we empirically find that increasing data capacity can prevent privacy leakage as utilizing the entire dataset shows much better privacy preservation than utilizing only the half, which implies that the model may have well-concealed privacy under proper data complexity. Since a lower-capacity model (considering the data complexity) can protect privacy better, the sparsity of the model can also be beneficial to privacy. Kaya et al. (2020) showed that regularization can mitigate some privacy risks while data augmentation techniques also help with privacy. The role of data augmentation was further studied and it was pointed out that only specific data augmentation techniques have such ability to mitigate privacy risks Kaya and Dumitras (2021); Yu et al. (2021). In addition, Yuan and Zhang (2022) found that traditional model pruning techniques do not work as well as the layer-wise architectural changes of the model in terms of reducing model capacities for privacy. Besides classification models, such privacy risk led by improper memorization also widely exists in models that are trained in various forms, e.g., regression learning Tarun et al. (2023) and self-supervised learning Wang et al. (2024).
@@
Figure 3: [1st row]: the distance to the decision boundary and MIAs accuracy; [2nd row]: the distance to the origin and the distance to the decision boundary. For a sample’s distance to the decision boundary, we use the difference between 1st and 2nd maximum prediction probabilities. The results are obtained from dozens of independent experiments. The blue charts ((a) & (c)) are from train set, and the green charts ((b) & (d)) are from test set. (ResNet18, CIFAR-100). This figure is excerpted from Fang and Kim (2024b).
@@ 4.2 Trade-Offs Between Privacy Preservation and Generalizability
Model-Level Solutions
@@
The model-level solutions aim to develop a mechanism to make the prediction distributions aligned on the model’s end. A classical model-level solution is differentially private stochastic gradient descent (DP-SGD) Abadi et al. (2016). It adds the noise into the optimizer to prevent the model from taking the (undesirable) easiest way to fit on the training data points and also memorizing them. Nasr et al. (2018) introduced an adversarial training framework (AdvReg) that mitigates membership inference attacks by aligning prediction distributions. It tries to develop a discriminator, similar to GAN, to identify the prediction inconsistency of the model while it makes the model try to deceive the discriminator to achieve prediction alignment. Chen et al. (2022) (RelaxLoss) established a threshold to prevent improper fitting for the alignment between member and non-member distributions while preserving the model’s generalizability through a technique similar to label smoothing. Tan et al. (2023b) proposed weighted smoothing (WS) to mitigate memorization by adding normalized random noise to the weights. Liu et al. (2024) incorporated a concave term called Convex-Concave Loss (CCLoss) to lessen the convexity of loss functions, aiming to enhance privacy preservation. Besides end-to-end solutions, there are also some studies exploring finer-grained solutions. Fang and Kim (2024b) introduced the Saturn Ring Classification Module (SRCM) to bound the representation magnitude to mitigate prediction disparity. Fang and Kim (2024a) tried to align representations in multiple layers by Center-based Relaxed Learning (CRL). Hu et al. (2024) proposed Privacy-Aware Sparsity Tuning (PAST) to measure weight-level privacy sensitivity and deactivate privacy-risky weights via regularization.
External Obfuscator
@@
The external obfuscator is a special kind of model-level solution. Instead of developing a privacy-safe model, it aims to build an obfuscator, which reproduces the prediction probabilities, to remedy the inconsistency in the prediction probabilities. Similar to the idea of DP-SGD, MemGuard Jia et al. (2019) interferes with the prediction confidence distribution of the model by adding additional noise after the model has been trained. Yang et al. (2023) tried to develop a VAE-based external prediction obfuscator named Purifier to align the prediction probabilities’ disparity. Different from MemGuard, it tries to reconstruct the prediction confidences to remove the prediction inconsistency instead of noise confusion.
Data-Level Solutions
@@
The data level approaches have two principles: training the privacy-safe model via (i) privacy-safe data or (ii) privacy-safe labels. It is straightforward that when all training data points are privacy-safe, there are no privacy-risky features included in the data, such as shortcut features Geirhos et al. (2020).
@@
The most straightforward solution to produce privacy-safer data is Data Augmentation. There are some data augmentation techniques, such as random cropping and flipping, determined that are able to produce privacy-safer data Kaya and Dumitras (2021); Yu et al. (2021). With augmented data, the model can usually achieve better privacy and generalizability. However, there are still no quantifiable metrics to measure how to further produce privacy-safe data through data augmentation. In other words, although the machine learning model can obtain privacy for free via data augmentation, it is unclear if the model achieves complete privacy safety yet. Stadler et al. (2022) tried to analyze the effect of synthetic data on the model’s privacy (Data Synthesis). Besides these two kinds of solutions, there is also an intuitive way to mitigate privacy risks. The first one is Data Pruning. With data pruning techniques, the model can use only a small amount of training data points to develop a well-generalized model. In other words, most membership privacy of the entire train set can still be protected. However, the privacy risk mitigation by data pruning is still not perfect, because some membership information of pruned data points could be leaked Li et al. (2024). The other one is Data Distillation. The data distillation aims to refine generalizability-critical features to produce some representative synthetic data. In this process, privacy-risky features can be removed Dong et al. (2022). As this direction has not been extensively researched yet, it is foreseen that more studies will be contributed to this topic very soon.
@@
Since the prediction disparity in training and testing is due to the improper fitting of the training data points, another way is to stop the model from further fitting into the data when it has learned enough information from the data. This means that if an ideal set of labels exists, the model can be trained perfectly privacy-safe on these labeled data. An intuitive idea is a distillation approach for membership privacy (DMP) Shejwalkar and Houmansadr (2021). It trains a protected model via non-member data and produces labels from an unprotected model. Another solution to better utilize limited data is self-ensemble architecture (SELENA) Tang et al. (2022). It developed an ensemble with an efficient sampling strategy to produce privacy-safe labels with better generalizability.

### v2

@@ 3.3 Limitations of Existing Techniques
All existing studies that aim to overcome spurious correlations work with the assumption that environments/groups that are overrepresented are the groups that contribute to the learning of spurious correlations. Reliance on this assumption makes it easy to identify which samples contain the spurious features causing problems, which allows for further representational alignment or changes to the training distribution. [44] show that this assumption does not always hold in practice and that minority groups can contain spurious features that can mislead a network significantly.
@@
Works by [38, 39] align representations of samples belonging to different environments within the same class. While effective at overcoming spurious correlations, these techniques reduce overall testing accuracies due to the loss of representational richness.
@@
Most works depend heavily on hyperparameter tuning, where they optimize for the best worst-group accuracy. Optimization is done with the help of a validation split that mimics the distribution of shifted testing environments. Such access to a validation split that mimics test-time distribution is unrealistic. [45] show that without access to such a validation set, standard Empirical Risk Minimization outperforms seemingly promising solutions.
@@ 3.4 Creating Robust Solutions: Next Steps
Most standard and state-of-the-art techniques assume an equal contribution to the learning and reliance of spurious correlations. In other words, every training sample belonging to the environment known to cause reliance on spurious correlations is treated the same way. [44] show that samples within an environment contribute differently to learning of spurious correlations and show that these differences are extreme in practice. They train a network to learn gender classification, where a fraction of the male samples contain eyeglasses. In their work, the degree of spurious feature reliance is measured by observing the test accuracy of female samples containing eyeglasses (Worst-Group Accuracy). They observe that removing 97% of easy-to-understand male samples with eyeglasses has almost no improvement on the testing accuracy of female samples with eyeglasses. However, removing 10% of hard-to-understand male samples with eyeglasses doubles the testing accuracy of female samples with eyeglasses, as shown in Fig. 1. They show that such pruning has no negative impact on overall testing accuracy.
@@ 3.5 Intersection of Spurious Correlations with Other Areas of Study
Recent work has shown that deep neural networks have a tendency to rely on short-cut solutions or heuristics when learning to solve reasoning tasks, instead of robust rules that actually cover the solution to the problem [46, 47]. This makes it difficult for networks to generalize to different or more challenging domains. A good example of this is the length generalization problem, where a network is unable to solve simple arithmetic operations on numbers of length different from those observed during training, despite these operations requiring the same set of rules [48, 49].
@@
[50] show that neural networks pick up on spurious features present in only a handful of training samples, which can lead to privacy leaks.
@@ 4.1 Privacy Correlation on Model Capacity and Data Complexity
The membership privacy risks of machine learning models are mainly caused by the model’s memorization of the training data points. This means that over-memorization is one of the sources of privacy risks [51]. [52] claimed some data points must be more privacy-risky after the removal of original privacy-risky data points and retraining from scratch. This is mainly due to the relative changes between the model capacity and the data complexity. [53] found that excess model capacity (a.k.a., overparameterization) is another factor of privacy risks. Additionally, [54] showed the larger-capacity model not only memorizes more on training data points than smaller networks but also memorizes faster (i.e., within fewer iterations). In fact, changing data complexity can also change the model’s memorization behavior. As shown in Fig. 2, we empirically find that increasing data capacity can prevent privacy leakage as utilizing the entire dataset shows much better privacy preservation than utilizing only the half, which implies that the model may have well-concealed privacy under proper data complexity. Since a lower-capacity model (considering the data complexity) can protect privacy better, the sparsity of the model can also be beneficial to privacy. [55] showed that regularization can mitigate some privacy risks while data augmentation techniques also help with privacy. The role of data augmentation was further studied and it was pointed out that only specific data augmentation techniques have such ability to mitigate privacy risks [56, 57]. In addition, [58] found that traditional model pruning techniques do not work as well as the layer-wise architectural changes of the model in terms of reducing model capacities for privacy. Besides classification models, such privacy risk led by improper memorization also widely exists in models that are trained in various forms, e.g., regression learning [59] and self-supervised learning [60].
@@
Figure 3: [1st row]: the distance to the decision boundary and MIAs accuracy; [2nd row]: the distance to the origin and the distance to the decision boundary. For a sample’s distance to the decision boundary, we use the difference between 1st and 2nd maximum prediction probabilities. The results are obtained from dozens of independent experiments. The blue charts ((a) & (c)) are from train set, and the green charts ((b) & (d)) are from test set. (ResNet18, CIFAR-100). This figure is excerpted from [61].
@@ 4.2 Trade-Offs Between Privacy Preservation and Generalizability
4.2.1 Model-Level Solutions
@@
The model-level solutions aim to develop a mechanism to make the prediction distributions aligned with the model’s end. A classical model-level solution is differentially private stochastic gradient descent (DP-SGD) [62]. It adds noise into the optimizer to prevent the model from taking the (undesirable) easiest way to fit the training data points and also memorizing them. [63] introduced an adversarial training framework (AdvReg) that mitigates membership inference attacks by aligning prediction distributions. It tries to develop a discriminator, similar to GAN, to identify the prediction inconsistency of the model while it makes the model try to deceive the discriminator to achieve prediction alignment. [64] (MixUp+MMD) further improved the defending ability by combining mixup data augmentation and maximum mean discrepancy based regularization. [65] (RelaxLoss) established a threshold to prevent improper fitting for the alignment between member and non-member distributions while preserving the model’s generalizability through a technique similar to label smoothing. [66] proposed weighted smoothing (WS) to mitigate memorization by adding normalized random noise to the weights. Combining the advantages of label-smoothing and MemGuard, [67] proposed HAMP, a privacy-beneficial solution on both training and inference stages. [68] incorporated a concave term called Convex-Concave Loss (CCLoss) to lessen the convexity of loss functions, aiming to enhance privacy preservation. Besides end-to-end solutions, there are also some studies exploring finer-grained solutions. [61] introduced the Saturn Ring Classification Module (SRCM) to bound the representation magnitude to mitigate prediction disparity. [69] tried to align representations in multiple layers by Center-based Relaxed Learning (CRL). [70] proposed Privacy-Aware Sparsity Tuning (PAST) to measure weight-level privacy sensitivity and deactivate privacy-risky weights via regularization. [71] (MIST) tried to remove the model’s privacy-risky bias via ensembling alignment.
4.2.2 External Obfuscator
@@
The external obfuscator is a special kind of model-level solution. Instead of developing a privacy-safe model, it aims to build an obfuscator, which reproduces the prediction probabilities, to remedy the inconsistency in the prediction probabilities. Similar to the idea of DP-SGD, MemGuard [72] interferes with the prediction confidence distribution of the model by adding additional noise after the model has been trained. [73] tried to develop a VAE-based external prediction obfuscator named Purifier to align the prediction probabilities’ disparity. Different from MemGuard, it tries to reconstruct the prediction confidences to remove the prediction inconsistency instead of noise confusion.
4.2.3 Data-Level Solutions
@@
The data level approaches have two principles: training the privacy-safe model via (i) privacy-safe data or (ii) privacy-safe labels. It is straightforward that when all training data points are privacy-safe, there are no privacy-risky features included in the data, such as shortcut features [31].
Privacy-Safe Data
@@
The most straightforward solution to produce privacy-safer data is Data Augmentation. There are some data augmentation techniques, such as random cropping and flipping, determined that are able to produce privacy-safer data [56, 57]. With augmented data, the model can usually achieve better privacy and generalizability. However, there are still no quantifiable metrics to measure how to further produce privacy-safe data through data augmentation. In other words, although the machine learning model can obtain privacy for free via data augmentation, it is unclear if the model achieves complete privacy safety yet. [74] tried to analyze the effect of synthetic data on the model’s privacy (Data Synthesis). Besides these two kinds of solutions, there is also an intuitive way to mitigate privacy risks. The first one is Data Pruning. With data pruning techniques, the model can use only a small amount of training data points to develop a well-generalized model. In other words, most membership privacy of the entire train set can still be protected. [75] (LOOD) is such a study identifying the privacy-risky data based on Leave-One-Out mechanism. However, the privacy risk mitigation by data pruning is still not perfect, because some membership information of pruned data points could be leaked [76]. The other one is Data Distillation. The data distillation aims to refine generalizability-critical features to produce some representative synthetic data. In this process, privacy-risky features can be removed [77]. As this direction has not been extensively researched yet, it is foreseen that more studies will be contributed to this topic very soon.
Privacy-Safe Labeling
@@
Since the prediction disparity in training and testing is due to the improper fitting of the training data points, another way is to stop the model from further fitting into the data when it has learned enough information from the data. This means that if an ideal set of labels exists, the model can be trained perfectly privacy-safe on these labeled data. An intuitive idea is a distillation approach for membership privacy (DMP) [78]. It trains a protected model via non-member data and produces labels from an unprotected model. Another solution to better utilize limited data is self-ensemble architecture (SELENA) [79]. It developed an ensemble with an efficient sampling strategy to produce privacy-safe labels with better generalizability.